OpenStudy (ria23):

Express the polynomial as a product of linear factors. f(x)=2x^3+4x^2-2x-4

OpenStudy (cwrw238):

factor in pairs first 2 and second 2

OpenStudy (ria23):

How do I know which is the first and the second...

OpenStudy (akashdeepdeb):

\[f(x)=2x^3+4x^2-2x-4\] Factorize! \[2x^3 + 4x^2 - 2x - 4 = 2x^2(x+2) - 2(x+2) = (2x^2 - 2)(x+2) = 2(x^2-1)(x+2) = 2(x+1)(x-1)(x+2)\]

OpenStudy (cwrw238):

Akash has done it for you...

OpenStudy (ria23):

:O Oh! Oh I get it... >u<

OpenStudy (cwrw238):

you can take the 2 out of first barackets so we have 2(x^2 - 1)(x + 2) now first bracket is difference of 2 squares;- = 2(x-1)(x+1)(x + 2)

OpenStudy (ria23):

Oh... I see it now.... c: I was confused on where the x+1 came from... I kind of get it now...

OpenStudy (ria23):

Hey cwcw, can I ask yhu to medal akash so both of yhu can get a medal? ;o

OpenStudy (akashdeepdeb):

Relax! I am not here to help you for medals. Understand the question and I am sure it'll please me.

OpenStudy (cwrw238):

yes - but i gave you a medal anyway

OpenStudy (ria23):

Haha, I know, most aren't... I just want yhu to know that I appreciate yhur help! Yhu and cwcw. Thank yhu guys for showing me how to do it. c: